OpenStudy (anonymous):

the velocity time graph of a ball of mass 25g moving on a road is as given below a)how much force does the road exert on the ball to bring it to rest. b)what is the direction of force exerted by the road.?

OpenStudy (anonymous):

13th question in the given QP

OpenStudy (ghazi):

Force= mass* aceleration acceleration = dV/dt= slope of velocity - time graph by this you will have retardation= 2.5m/s^2 therefore force= 0.025*2.5=0.0625N

OpenStudy (ghazi):

now direction of force exerted on ball will be opposite to the direction of motion of ball

OpenStudy (anonymous):

F=mass * acceleration Mass is given = 25 g = 0.025 kg Acceleration would be the gradient of the graph since it is a velocity vs. time graph and m/s/t = m/s/s = the units for acceleration.

OpenStudy (ghazi):

acceleration=(0-10)/(4-0)=-2.5m/s^2 that is retardation